Grade 12Algebra

Multiply and reduce to lowest form and convert into a mixed fraction:
11 × (4/7)

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

By the rule Multiplication of fraction,
Product of fraction = (product of numerator)/ (product of denominator)
Then,
= (11/1) × (4/7)
= (11 × 4)/ (1 × 7)
= (44/7)
= 6 whole 2/7

Thanks